OpenStudy (anonymous):

simplify the trigonometric expression? sin^2(theta)/1- cos(theta)

Nnesha (nnesha):

trig identity sin^2 theta = what ?

OpenStudy (anonymous):

Just a suggestion, but go to mathway.com and it solves most problems your stuck on and helps break stuff down

OpenStudy (anonymous):

i know sin (theta) = 1/csc(theta)

Nnesha (nnesha):

just becareful sometimes it will give u the wrong answer js :)

Nnesha (nnesha):

identity \[\huge\rm sin^2 \theta + \cos^2 \theta = 1 \] solve this for sin^2 theta

OpenStudy (anonymous):

oh yeah, i forgot about that! sin^2 (theta) = 1-cos^2(theta)

OpenStudy (anonymous):

Yeah you have to be exact with it and sometimes it wont give you the answer so, Yeah:) very helpful though still

Nnesha (nnesha):

yep right now apply difference of square method \[\huge\rm a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

(1 + cos (theta))(1-cos(theta))?

OpenStudy (anonymous):

so the answer is 1 + cos(theta)?

Nnesha (nnesha):

yep right \[\huge\rm \frac{ (1 - \cos \theta )(1+ \cos \theta)}{ 1-\cos \theta }\]

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

cuz 1 - cos(theta) cancels out

OpenStudy (anonymous):

thanks!

Nnesha (nnesha):

np :-)